OpenStudy (anonymous):

Geometry help please? In an isosceles triangle ABC, AB = AC A = 20* Point P is taken on side AC such that AP = BC. Find PBC.

OpenStudy (badhi):

$$BC=2AC\sin(10^\circ)$$ $$PC=AC-2AC\sin(10^\circ)=AC[1-2\sin(10^\circ)]$$ from sine rule to PCB triangle, $$\frac{PC}{\sin \alpha}=\frac{BC}{\sin[180-\alpha-80]}$$ now solve for alpha

OpenStudy (anonymous):

u know it then plz tell it

OpenStudy (anonymous):

why so much suspense

OpenStudy (anonymous):

in triangle abc ab=ac ap=bc <a=20 to find <pbc

OpenStudy (anonymous):

after this what should i do

OpenStudy (anonymous):

and i didnt understan what badhi told

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so u have any solution

OpenStudy (anonymous):

no i dont have the answer

OpenStudy (anonymous):

i never said that

OpenStudy (anonymous):

oh then pls tell

OpenStudy (anonymous):

i forgot to put the q mark

OpenStudy (anonymous):

have any idea?

OpenStudy (anonymous):

do we have to use trigo here

OpenStudy (anonymous):

rotation

OpenStudy (anonymous):

any other method that we can use here

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

mertsj

OpenStudy (anonymous):

have any ideas

OpenStudy (mertsj):

Do you want to talk or do you want to work on this problem?

OpenStudy (anonymous):

no uh actually i have to go now so we will meet tomorrow same time sama place

OpenStudy (anonymous):

please tell me the sol after wards

OpenStudy (mertsj):

Bye bye

OpenStudy (anonymous):

will meet tomorrow?

OpenStudy (mertsj):

We'll see. Maybe later today. I'll post a solution if I find one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

still not got

OpenStudy (anonymous):

is that the angle of A you have given or a length??

OpenStudy (anonymous):

angle

OpenStudy (anonymous):

HOW CAN ANGLE B AND ANGLE C BE 170

OpenStudy (anonymous):

cosine rule \[A ^{2}=B ^{2}+C ^{2}-2BCcosA ^{*}\]

OpenStudy (anonymous):

BUT ANGLE B AND ANGLE C ARE NOT 170

OpenStudy (anonymous):

AS ANGLE IN A TRIANGLE IS 180

OpenStudy (anonymous):

i am sorry it is rather